Calculate the weight (kg) of parboiled with 40% MC (wb) required yield 1 ton of desired paddy with 12% MC (wb)
1
Expert's answer
2020-07-15T18:30:42-0400
"Solution"

"W_f=W_i*\frac{100-MC_i}{100-MC_f}"


"MC_{i}=40" %


"MC_{f}=12" %


"MC_{wb}=" "Moisture" "Content" "Wet" "Basis" [%]


"W_i=" "Initial" "Weight"


"W_f=" "Final" "Weight=" "1 ton" "=1000kg"


"1000kg=W_i(\frac{100-40}{100-12})=W_i(\frac{60}{88})"


"W_i=\frac{1000}{\frac{60}{88}}=1466.70577kg" "of" "paddy" "at" "40"% "MC_{wb}"